how do i graph x+2y=-2 useing intercept method?
Found 2 solutions by solver91311, rfer:
Answer by solver91311(24713)



Substitute 0 for and then do the arithmetic to solve for . Plot the point using the value you just calculated as the -coordinate.

Substitute 0 for and then do the arithmetic to solve for . Plot the point using the value you just calculated as the -coordinate.

Draw a line through the two points you just plotted.

John


Answer by rfer(16322)

y int when x=0 (0,-1)
x int when y=0 (-2.0)
------------
x+2y=-2
0+2y=-2
y=-1
------------
x+2(0)=-2
x=-2